OpenStudy (anonymous):

you have 20 y ball, 55 g balls, 25 p balls. you pick one ball at random. what is the theoretical probabilty of picling a p ball

OpenStudy (anonymous):

\[\text{number of p balls} \over \text{total number of balls}\]

OpenStudy (anonymous):

exactly for Y ball probability is 20/100=1/5 for g ball probability is 55/100 for p ball probability is 25/100=1/4